# Description of the structure of DG–algebras via characters on a groupoid

### v2

@@ Abstract
We construct a description of graded derivations in group algebras. Using this result for arbitrary graduation of the group algebra, we describe all possible structures of DG algebras. The corresponding examples are given. The description is given in terms of characters on a groupoid analogous to the groupoid of inner action.
@@ Introduction
In several previous papers one offers A.S. Mishchenko and others [AMS17, AM19, AAS23], was presented a method for studying the derivations of group algebras using the character space on the groupoid of conjugacy action. In the present paper, a similar groupoid and characters on it are constructed in order to describe graded derivations on graded group algebra and all possible structures of DG algebras on group algebras. [^1] The results of Sections 1.1 and 1.2 were obtained by A. Arutyunov, the other results were obtained jointly. The work of A. Arutyunov was supported by the RSF grant No. 25-11-00018.
DG algebra and graded derivations are frequently encountered. See e.g. [Orl23] where they are studied in the context of algebraic geometry. The independent study of graded derivations is an even better known question, see e.g. [Mei03]. The classification of possible DG strings has also been the focus of study [Mao11, MWZ22].
The main object of our study will be graded group algebra $\mathbb{C}[G]\cong\bigoplus A_{i}$ and its graded derivation. That is, such linear maps $d:\mathbb{C}[G]\rightarrow\mathbb{C}[G]$ for which holds the graded Leibniz rule
@@
In Section 1 we present a structure of a grouppod which is connected to the graded algevra structure. We will give a description of graded derivations as localy finite characters on this grouppoid (see Theorem 1.1) In Section 2, we describe the inner, quasi-inner, and central derivations for the graded case. Using this construction in the case of DG–algebras, in Section 3 we obtain a description of the structure of the DG–algebra in terms of characters and an isomorphism criterion between two DG–algebras. In addition, at the end of Section 3, we give several specific examples of graduations on group algebras and derivations that define the structure of the DG–algebra on them. The definitions and properties from combinatorial group theory that are necessary for the paper can be found in [LS01]
1 Groupoid, characters and graded derivations.
Let’s start with the definitions and fix the notation for future reference. We will study derivations over group algebras with a DG structure. A group algebra is defined in standard form as follows:
@@
$K[G]$ – is the group algebra of a group $G$ over a field $K$ is an associative algebra over a field $K$, the elements of which are all possible formal finite sums of the form $\sum_{g\in G}k_{g}g,$ $g\in G,$ $k_{g}\in K$, and the operations are determined by the formulas:
@@
$(\sum_{g\in G}a_{g}g)(\sum_{g\in G}b_{g}g)=\sum_{h\in G}(\sum_{xy=h\in G}(a_{x}b_{y})h).$
The elements of the group $G$ form the basis of the algebra $K[G]$. The multiplication of the basic elements in the group algebra is induced by the multiplication in group.
Furthermore, we need the concept of a groupoid. Typically, this is a category in which every morphisms is invertible. More precisely, we will work with the following construction associated with a finitely generated group $G$, so this category is small (objects and morphisms are sets).
@@
A groupoid $\Gamma$ contains a set of objects $Obj(\Gamma)$ and morphisms $Hom(\Gamma)$. Each morphism has a source $s(\phi)$ and a target $t(\phi)$, i.e. each morphism is in the set $Hom(s(\phi),t(\phi))$. Composable morphisms have an associative operation (composition). The following conditions hold:
@@
    For each object $a\in Obj(\Gamma)$, there is a neutral endomorphism $1_{a}\in Hom(a,a)$ such that $\forall\phi\in Hom(b,a)$ $1_{a}\circ\phi=\phi$ and $\forall\phi\in Hom(a,b)$ $\phi\circ 1_{a}=\phi$.
@@
    There is an inverse for each morphism. I.e. $\forall\phi\in Hom(a,b)$ exists $\psi\in Hom(b,a)$ such that $\phi\circ\psi=1_{b}$ and $\psi\circ\phi=1_{a}$.
@@
The following construction is a generalization of the structure from [AMS17]. Suppose we have a graded group algebra $\mathbb{C}[G]\cong\bigoplus A_{i}$, here $i\in\mathbb{Z}$. Group $G$ assumes to be finitely generated. We will define the groupoid of conjugacy action, denoted by $\Gamma$, for the group $G$ in the sense of Definition 1.2.
@@
Consider two morphisms $\phi=(u_{1},v_{1})$ and $\psi=(u_{2},v_{2})$ such that $t(\phi)=s(\psi)$, that is, the composition is defined for them $\psi\circ\phi$. Define the composition of morphisms as follows
$(u_{2},v_{2})\circ(u_{1},v_{1}):=((-1)^{|v_{2}|}v_{2}u_{1},v_{2}v_{1})).$
This formula is conveniently recorded in the form of a diagram.
@@
Let we have $\phi=(u_{1},v_{1})$ and $\psi=(u_{2},v_{2})$ such that $t(\phi)=s(\psi)$. Let’s say $s(\phi)=a$, $t(\phi)=s(\psi)=b$ and $t(\psi)=c$.
Check that
@@
By definition
$s(\psi\circ\phi)=(-1)^{|v_{2}v_{1}|}v_{1}^{-1}v_{2}^{-1}(-1)^{|v_{2}|}v_{2}u_{1}$ = $(-1)^{|v_{1}|}v_{1}^{-1}u_{1}=a$.
@@
$t(\psi\circ\phi)$ = $(-1)^{|v_{2}|}v_{2}u_{1}v_{1}^{-1}v_{2}^{-1}=u_{2}v_{2}^{-1}=c$, because $u_{1}v_{1}^{-1}=(-1)^{|v_{2}|}v_{2}^{-1}u_{2}$
Now we prove the associativity of the composition. Let $\phi_{1},\phi_{2},\phi_{3}$ - morphisms such that $s(\phi_{1})=t(\phi_{2})$ and $s(\phi_{2})=t(\phi_{3})$, then
@@
We proved above that all compositions are defined correctly. Let’s calculate the left part.
$\phi_{2}\circ\phi_{3}=((-1)^{|v_{2}|}v_{2}u_{3},v_{2}v_{3}))$ $\phi_{1}\circ(\phi_{2}\circ\phi_{3})=(u_{1},v_{1})\circ((-1)^{|v_{2}|}v_{2}u_{3},v_{2}v_{3}))=((-1)^{|v_{1}|+|v_{2}|}v_{1}v_{2}u_{3},v_{1}v_{2}v_{3}).$
Now let’s calculate the right side of the expression
$(\phi_{1}\circ\phi_{2})\circ\phi_{3}=((-1)^{|v_{1}|}v_{1}u_{2},v_{1}v_{2})\circ(u_{3},v_{3})=((-1)^{|v_{1}|+|v_{2}|}v_{1}v_{2}u_{3},v_{1}v_{2}v_{3}).$
The left and right sides of the expression are equal, which means that the operation $\circ$ is associative.
Let’s show that there is an element $1_{g}$, such that:
$s(1_{g})=t(1_{g})=g$
$\phi\cdot 1_{s(\phi)}=\phi=1_{t(\phi)}\cdot\phi$, $\forall\phi\in$ $Hom(\Gamma)$
Let’s take as a $1_{g}$ element $(g,e)$, where e is a trivial element of the group G.
@@
Indeed, let’s take an arbitrary morphism $(u,v)$, than:
$(uv^{-1},e)(u,v)=(u,v)$.
@@
We have shown that $\Gamma$ is a category. Now it is necessary to check the inversibility of the arrows.
We show the presence of an inverse morphism $\forall\phi\in Hom(\Gamma)$. For morphism of the form (u,v) let’s call the inverse morphism $((-1)^{|v|}v^{-1}uv^{-1},v^{-1})$.
Indeed:
$(u,v)((-1)^{|v|}v^{-1}uv^{-1},v^{-1})=(uv^{-1},e)=(t(\phi),e).$
@@
Thus, the Proposition 1.1. is proved. ∎
@@
    For endomorphisms $Hom(a,a),a\in\pm G$ holds the following descriptions
    $Hom(a,a)=\{\chi\in Hom(\Gamma)\mid s(\chi)=(-1)^{|t|}t^{-1}a,$ $t(\chi)=at^{-1},$ $t\in Z(a)$ and $t\in$ $A_{2i}[G]\}$
@@
    For $a]:=\{b\in\pm G\mid\exists t\in\pm G:b=(-1)^{|t|}t^{-1}at\}$ denote by $\Gamma_{[a]}$, a subgroupoid such that $Obj(\Gamma_{[a]})=[a]$. The following properties holds:
@@
      $Hom(\Gamma_{[a]})=\{\phi\in Hom(\Gamma)\mid s(\phi)\in[a],t(\phi)\in[a]\}$.
@@
    There is a left action $Hom(a,a)\times Hom(a,b)\rightarrow Hom(a,b)$, $(\phi,\psi)\to\psi\circ\phi$. This action is free and transitive.
Let us show that with the help of characters of $\Gamma$, it is possible to describe a derivation satisfying the graded Leibniz rule.
@@
We say that the linear operator $d$: $\mathbb{C}[G]$ $\to$ $\mathbb{C}[G]$ satisfies the graded Leibniz rule if
@@
We call the map $\chi$ : $Hom(\Gamma)$ $\to$ $\mathbb{C}$ a character if it satisfies the identity $\chi(\psi\circ\phi)=\chi(\psi)+\chi(\phi)$
@@
We will say that the character $\chi$ is locally finite if for all $v\in pmG$ we get that $\chi(u,v)=0$ for all except a finite number $u$.
We will denote by $\chi(\Gamma)$ the space of local finite characters. In section 1.2 we will show that $\chi(\Gamma)$ has a natural structure of Lie algebra.
The graded Leibniz rule in a group algebra can be written in terms of characters on the groupoid $\Gamma$ as follows.
@@
For any graded derivation $d$ there exists a locally finite character such that
$d(u)=\sum\limits_{g\in G}\Big{(}\sum\limits_{h\in\pm G}\chi(\phi(h,g)\lambda^{h}\Big{)}g,\forall u\in A_{i}$ (1)
@@
It is enough to check the statement on the elements $g\in\pm G$, since they generate $\mathbb{C}[G]$, as a vector space.
For any element $g$ of group G we have $d(g)=\sum\limits_{h\in G}d_{g}^{h}h$
Define $\chi(\phi(h,g)):=d_{g}^{h}$, $\chi(\phi(-h,g)):=-d_{g}^{h}$ and let us show that it is a character. We have
$d(g_{2}g_{1})=d(g_{2})g_{1}+(-1)^{|g_{2}|}g_{2}d(g_{1}).$
Let’s rewrite this in a new notation:
@@
It is clear that summation over all $h$ from the group for $h$ and $gh$ is equivalent, so we can convert the right-hand sides to the following form:
@@
A corollary of the above statement is the following theorem, the proof of which is similar to the non-graded case, see Theorem 1 of [Aru20].
@@
For any graded derivation $d$ there exists such a unique locally finite character of $\chi_{d}$, that the formula is valid
@@
In fact, we can say that we have an isomorphism $Der(\mathbb{C}[G])\cong\chi(\Gamma)$.
To prove this, we need to use formula (1) to element $a\in\pm G$ and write an arbitrary element $g\in G$ as $g=at$ and then remove the terms with a zero coefficient from the sum. See the proof of Theorem 1 in [Aru20] for more details.
@@
Locally finite characters are isomorphic to derivations, which means that the binary commutator operation on derivations generates a binary operation on characters. Formula (1) allows us to define a composition in character space. Let $d_{1},d_{2}$ be the derivations, defined by the characters $\chi_{d_{1}},\chi_{d_{2}}$. The character that corresponds to the commutator of derivations $d_{1},d_{2}]$ is denoted by $\chi_{[d_{1},d_{2}]}$. It defines an operation on the character space.
$\{\chi_{d_{1}},\chi_{d_{2}}\}:=\chi_{[d_{1},d_{2}]}$ (3)
For this operation, we have the following equality
@@
Holds the formula
@@
Due to local finiteness we can change the order of summation in the last equation and get
@@
It follows that the pair $(\chi_{\Gamma},\{\cdot;\cdot\})$ is a Lie algebra which is isomorphic by proposition 1.3 to the algebra of all graded derivations.
2 Examples of graded derivations and their description using characters on a groupoid.
We will demonstrate the correspondence between characters on groupoid and graded derivations using examples of inner derivations, central derivations and quasi-inner derivations. It gives give easy examples of such operators.

### v3

# Description of DG–algebras structure via characters on a groupoid

@@ Abstract
We construct a description of graded derivations in group algebras. Using this result for arbitrary grading of the group algebra, we describe all possible DG–algebras structures. Examples are given. The description is given in terms of characters on a groupoid analogous to the groupoid of conjugate action.
@@ Introduction
In several previous papers bye A.S. Mishchenko and others [AMS17, AM19, AAS23], was presented a method for studying the derivations of group algebras using the character space on the groupoid of conjugacy action. In the present paper, a similar groupoid and characters on it are constructed in order to describe graded derivations on graded group algebras and all possible DG–algebra structures on group algebras. [^1] The results of Sections 1.1 and 1.2 were obtained by A. Arutyunov, the other results were obtained jointly. The work of A. Arutyunov was supported by the RSF grant No. 25-11-00018.
DG–algebras and graded derivations are occur frequently. See e.g. [Orl23], where they are studied in the context of algebraic geometry. Graded derivations in isolation are an even better known subject, see e.g. [Mei03]. The classification of possible DG structures has also been a focus of study [Mao11, MWZ22].
The main object of our study will be the graded group algebra $\mathbb{C}[G]\cong\bigoplus A_{i}$ and its graded derivations, that is linear maps $d:\mathbb{C}[G]\rightarrow\mathbb{C}[G]$ satisfying the graded Leibniz rule
@@
In Section 1 we present the structure of a groupoid which is connected to the graded algebra structure. We will give a description of graded derivations as locally finite characters on this groupoid (see Theorem 1.1). In Section 2, we describe the inner, quasi-inner and central derivations in the graded case. Using this construction in the case of DG–algebras, in Section 3 we obtain a description of DG–algebra structures in terms of characters and an isomorphism criterion for DG–algebras. In addition, at the end of Section 3, we give several examples of gradings on group algebras and derivations that define a DG–algebra structure on them. The definitions and properties from combinatorial group theory that are necessary to understand the present paper can be found in [LS01]
1 The groupoid, it’s characters, and graded derivations.
Let’s start with the definitions and fix the notation for future reference. We will study derivations of group algebras with a DG structure. The standard definition of a group algebra is
@@
Let $G$ be a group and let $K$ a field. The group algebra $K[G]$ is the associative algebra over $K$ the elements of which are all possible formal finite sums of the form $\sum_{g\in G}k_{g}g,$ $g\in G,$ $k_{g}\in K$, and the operations are defined by:
@@
$(\sum_{g\in G}a_{g}g)(\sum_{g\in G}b_{g}g)=\sum_{h\in G}\Big{(}\sum_{xy=h\in G}(a_{x}b_{y})h\Big{)}.$
The elements of the group $G$ form a basis of $K[G]$. Multiplication of the basis elements in the group algebra is induced by the group multiplication.
Furthermore, we need the notion of a groupoid. usually, it is defined as a category in which every morphism is invertible. We will work with the following construction associated with a finitely generated group $G$, so this category is small (the classes of objects and morphisms are sets).
@@
A groupoid $\Gamma$ has a set of objects $\operatorname{Obj}(\Gamma)$ and a set of morphisms $\operatorname{Hom}(\Gamma)$. Each morphism has a source $s(\phi)$ and a target $t(\phi)$, i.e. each morphism is in the set $\operatorname{Hom}(s(\phi),t(\phi))$. Composable morphisms have an associative operation (composition). The following conditions hold:
@@
    For each object $a\in\operatorname{Obj}(\Gamma)$, there is a neutral endomorphism $1_{a}\in\operatorname{Hom}(a,a)$ such that $\forall\phi\in\operatorname{Hom}(b,a)$: $1_{a}\circ\phi=\phi$ and $\forall\phi\in\operatorname{Hom}(a,b)$ $\phi\circ 1_{a}=\phi$.
@@
    Each morphism has the inverse: fir every $\phi\in\operatorname{Hom}(a,b)$ there exists $\psi\in\operatorname{Hom}(b,a)$ such that $\phi\circ\psi=1_{b}$ and $\psi\circ\phi=1_{a}$.
@@
The following construction generalizes the one from [AMS17]. Suppose we have a $\mathbb{C}$–graded group algebra $\mathbb{C}[G]\cong\bigoplus A_{i}$. The group $G$ is assume to be finitely generated. We will define the groupoid of conjugacy action $\Gamma$ for the group $G$ (in the sense of Definition 1.2) as follows:
@@
Consider two morphisms $\phi=(u_{1},v_{1})$ and $\psi=(u_{2},v_{2})$ such that $t(\phi)=s(\psi)$, that is, the composition $\psi\circ\phi$ should be defined. Define it as follows:
$(u_{2},v_{2})\circ(u_{1},v_{1}):=((-1)^{|v_{2}|}v_{2}u_{1},v_{2}v_{1}).$
This formula is conveniently depicted as a diagram.
@@
Consider $\phi=(u_{1},v_{1})$ and $\psi=(u_{2},v_{2})$ such that $t(\phi)=s(\psi)$. Define $s(\phi)=a$, $t(\phi)=s(\psi)=b$ and $t(\psi)=c$.
We need to check that
@@
By definition,
$s(\psi\circ\phi)=(-1)^{|v_{2}v_{1}|\mathstrut}v_{1}^{-1\mathstrut}v_{2}^{-1\mathstrut}(-1)^{|v_{2}|\mathstrut}v_{2}u_{1}$ = $(-1)^{|v_{1}|\mathstrut}v_{1}^{-1\mathstrut}u_{1}=a$.
@@
$t(\psi\circ\phi)$ = $(-1)^{|v_{2}|\mathstrut}v_{2}u_{1}v_{1}^{-1\mathstrut}v_{2}^{-1\mathstrut}=u_{2}v_{2}^{-1\mathstrut}=c$, because $u_{1}v_{1}^{-1\mathstrut}=(-1)^{|v_{2}|\mathstrut}v_{2}^{-1\mathstrut}u_{2}$
Here we shall prove that the composition is associative. Let $\phi_{1},\phi_{2},\phi_{3}$ be morphisms such that $s(\phi_{1})=t(\phi_{2})$ and $s(\phi_{2})=t(\phi_{3})$. We need to show that
@@
We have proved above that these compositions are well–defined. The left side is
$\phi_{2}\circ\phi_{3}=((-1)^{|v_{2}|}v_{2}u_{3},v_{2}v_{3}))$ $\phi_{1}\circ(\phi_{2}\circ\phi_{3})=(u_{1},v_{1})\circ((-1)^{|v_{2}|}v_{2}u_{3},v_{2}v_{3}))=((-1)^{|v_{1}|+|v_{2}|}v_{1}v_{2}u_{3},v_{1}v_{2}v_{3}),$
While the right one is equal to it,
$(\phi_{1}\circ\phi_{2})\circ\phi_{3}=((-1)^{|v_{1}|}v_{1}u_{2},v_{1}v_{2})\circ(u_{3},v_{3})=((-1)^{|v_{1}|+|v_{2}|}v_{1}v_{2}u_{3},v_{1}v_{2}v_{3}),$
proving associativity.
Let us show that there is an element $1_{g}$ such that
$s(1_{g})=t(1_{g})=g$,
$\phi\cdot 1_{s(\phi)}=\phi=1_{t(\phi)}\cdot\phi$, $\forall\phi\in$ $\operatorname{Hom}(\Gamma).$
Take $1_{g}=(g,e)$, where $e$ is the identity of the group $G$.
@@
Indeed, choosing an arbitrary morphism $(u,v)$, we have:
$(uv^{-1},e)(u,v)=(u,v)$;
@@
We have now shown that $\Gamma$ is a category. It remains to check the invertibility of the arrows.
We claim the the inverse of a morphism of the form $(u,v)$ is $((-1)^{|v|}v^{-1}uv^{-1},v^{-1})$.
Indeed,
$(u,v)((-1)^{|v|}v^{-1}uv^{-1},v^{-1})=(uv^{-1},e);=(t(\phi),e).$
@@
This proves Proposition 1.1. ∎
@@
    The endomorphisms $\operatorname{Hom}(a,a),a\in\pm G$, admit the following description:
    $\operatorname{Hom}(a,a)=\{\chi\in\operatorname{Hom}(\Gamma)\mid s(\chi)=(-1)^{|t|}t^{-1}a,$ $t(\chi)=at^{-1},$ $t\in Z(a)$ and $t\in$ $A_{2i}[G]\}$
@@
    For $a]:=\{b\in\pm G\mid\exists t\in\pm G:b=(-1)^{|t|}t^{-1}at\}$, denote by $\Gamma_{[a]}$ the subgroupoid such that $\operatorname{Obj}(\Gamma_{[a]})=[a]$. The following properties hold:
@@
      $\operatorname{Hom}(\Gamma_{[a]})=\{\phi\in\operatorname{Hom}(\Gamma)\mid s(\phi)\in[a],t(\phi)\in[a]\}$.
@@
    The left action $\operatorname{Hom}(a,a)\times\operatorname{Hom}(a,b)\rightarrow\operatorname{Hom}(a,b)$: $(\phi,\psi)\to\psi\circ\phi$ is free and transitive.
Let us show that, with the help of characters of $\Gamma$, it is possible to give a description of derivations satisfying the graded Leibniz rule.
@@
The linear operator $d$: $\mathbb{C}[G]$ $\to$ $\mathbb{C}[G]$ satisfies the graded Leibniz rule if
@@
The map $\chi$ : $\operatorname{Hom}(\Gamma)$ $\to$ $\mathbb{C}$ is a character if it satisfies the identity $\chi(\psi\circ\phi)=\chi(\psi)+\chi(\phi)$:
@@
A character $\chi$ is locally finite if for all $v\in\pm G$ we have $\chi(u,v)=0$ for all except a finite number of $u$.
We will denote by $\chi(\Gamma)$ the space of locally finite characters. In section 1.2 we will show that $\chi(\Gamma)$ has a natural Lie algebra structure.
The graded Leibniz rule in a group algebra can be expressed in terms of characters on the groupoid $\Gamma$ as follows.
@@
For any graded derivation $d$ there exists a locally finite character $\chi$ such that
$d(u)=\sum\limits_{g\in G}\Big{(}\sum\limits_{h\in\pm G}\chi(\phi(h,g)\lambda^{h}\Big{)}g;\forall u\in A_{i}.$ (1)
@@
It is enough to check the statement on the elements $g\in\pm G$, since they generate $\mathbb{C}[G]$ as a vector space.
For any element $g$ of group G we have $d(g)=\sum\limits_{h\in G}d_{g}^{h}h$.
Let us define $\chi(\phi(h,g)):=d_{g}^{h}$, $\chi(\phi(-h,g)):=-d_{g}^{h}$ and show that $\chi$ it is a character. We have
$d(g_{2}g_{1})=d(g_{2})g_{1}+(-1)^{|g_{2}|}g_{2}d(g_{1}),$
or after a basis expansion
@@
It is clear that summing over all $h$ and all $gh$ is equivalent, so we can convert the right-hand side to the following form:
@@
A corollary of the above statement is the following theorem, proved similarly to the non-graded case, see Theorem 1 of [Aru20].
@@
For any graded derivation $d$ there exists a unique locally finite character $\chi_{d}$ satisfying
@@
There is therefore an isomorphism $\operatorname{Der}(\mathbb{C}[G])\cong\chi(\Gamma)$.
Proof.
To prove this, we need to apply formula (1) to $a\in\pm G$, write an arbitrary element $g\in G$ as $g=at$ and then drop the terms with a zero coefficient from the sum. See the proof of Theorem 1 in [Aru20] for more details. ∎
@@
Locally finite characters are isomorphic to derivations, which means that the commutator of derivations yields a binary operation on characters. Formula (1) allows us to define a commutator in character space. Let $d_{1},d_{2}$ be the derivations defined by the characters $\chi_{d_{1}},\chi_{d_{2}}$. Denote the character that corresponding to the commutator $d_{1},d_{2}]$ by $\chi_{[d_{1},d_{2}]}$. This defines an operation on the space of characters.
$\{\chi_{d_{1}},\chi_{d_{2}}\}:=\chi_{[d_{1},d_{2}]}.$ (3)
For this operation, we have the following equality:
@@
The operation on characters above can be expressed as
@@
Thanks to local finiteness, we can swap the sums in the last equation and get
@@
It follows that the $(\chi_{\Gamma},\{\cdot;\cdot\})$ is a Lie algebra, isomorphic to the algebra of all graded derivations by proposition 1.3.
2 Examples of graded derivations and their description in terms of characters.
We will illustrate the correspondence between characters on a groupoid and graded derivations using examples of inner derivations, central derivations and quasi-inner derivations. They will serve as easy examples of such operators.